OpenStudy (anonymous):

36^(x-2)=6 solve x show work

OpenStudy (anonymous):

Take the logarithm of both sides and isolate x

OpenStudy (anonymous):

i got it is 2.5

OpenStudy (anonymous):

correct

ganeshie8 (ganeshie8):

\(\large 36^{x-2}=6 \) \(\large (6^2)^{x-2}=6^1 \) \(\large 6^{2x-4}=6^1 \) since bases are equal, u can equatte the exponents \(\large 2x-4 = 1\) solve x

ganeshie8 (ganeshie8):

thats another way.. gives u same answer :)